

********************************************************************
UARS WEEKLY STATUS REPORT - 20 March 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM resumed operations on March 17. Some communication problems
were encountered in the KCRT/MSOCC system during the turn on.
Since then, SUSIM successfully gathered its normal solar,
offset, and occultation scan data without apparent problems.
***********
PEM
***********
PEM is currently off.
PEM is awaiting evaluation by GSFC of its proposed activation procedure.
The minimal configuration for PEM AUX activation is the zenith VMAG, MEPS1,
MEPS2, and HEPS1 units, only. This configuration is driven by the capabilities
of the spacecraft OBC and may increase if OBC program area increases. PEM
activation is proposed to be OBC driven when there exists plenty of power to
operate PEM. It is anticipated that the OBC will activate PEM on the daylight
portions of every orbit.
PEM supported the UARS Science team meeting this March 16. The PEM PI,
J. D. Winningham, gave a status update on PEM. PEM gave two Oral presentations
on Monday, March 16: "The UARS PEM Energetic Particle Climatology - Status and
Applications," (J. R. Sharber, R. A. Frahm, G. Crowley, A. J. Ridley, R. Link,
J. D. Winningham, E. E. Gaines, R. W. Nightingale, D. L. Chenette, and
W. R. Keith) and "Particle Effects on the Middle Atmosphere Modeled using the
Parallelized TIME-GCM," (G. Crowley, A. J. Ridley, J. D. Winningham, R. Link,
R. A. Frahm, J. R. Sharber, R. G. Roble, A. Richmond, and J. M. Russell III);
and a poster on Tuesday, March 17: "A Global PEM Electron Precipitation
Climatology for the Kp Index: Work in Progress," (R. A. Frahm, G. Crowley,
J. R. Sharber, J. D. Winningham, R. Link, A. J. Ridley, E. E. Gaines,
D. L. Chenette, B. J. Anderson and T. A. Potemra). All these PEM presentations
dealt with creating a PEM particle based climatology.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
The MLS instrument has been powered off since Mar 16. MLS is scheduled to
resume south viewing operations on Mar 26.
***********
HRDI
***********
HRDI is on and operating normally.
The HRDI rack node SPRLJ is down. Mail may be sent to HRDID instead of SPRLJ.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF worked with MPG, CMS and RAC Instrumentors to facilitate
communications directly with MPG and CMS in Building 32, without routing
through Building 23. Routers in B-23 were temporarily shut down Tuesday as a
test to ensure that PI's could reach MPG and CMS without the use of CDHF
routers. This activity is being done in the near term for the B-23 power
outage scheduled March 20-23.
The CDHF has received a total of 379 CD-ROM requests as of 20-March-1998.
There were two orders received and shipped.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
25 days PEM (3.32) PROD (1379R-1029C=350L)
122 days MLS (4.90) PROD (1297R-1055C=226L)
4 days CLAES (7.0) ENG (38R-23C=15L)
2 days HAL (4.80) ENG (4R-2C=2L)
Pending software upgrades: IDL 5.0, OpenIngres, UCSS.
***********
FOT
***********
REPORTING PERIOD
This report covers 06 March 1998 (Orbit 35435, GMT Day 066) through 13
March 1998 (Orbit 35539, GMT Day 072).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
SUSIM 070/2131 35508 Off
ACRIM 070/2311 35509 Off
HRDI 070/2311 35509 Off
MLS 070/2311 35509 On
WINDII 072/0255 35527 On
HALOE 072/1958 35537 Off
HALOE 072/2319 35539 On
The beta angle is now increasing toward a maximum angle of 54.5 degrees on
15 March 1998 (DOY 074).
Relative Time Sequence (RTS) number 12 , RTS-12, was generated and uplinked
to support cycling of compensation heater 18 (CH18) during the period of
time when SUSIM is powered off. The RTS was uplinked on 11 March 1998
during the 070/21:31 event - the same event in which SUSIM was powered off.
Through PMON initiated commanding, the RTS will enable both the A and B
sides of compensation heater 18, in an effort to keep the SUSIM instrument
warm during the power OFF phase. The original version of RTS-12 (which
enables only the A side of CH18) will be uplinked again when SUSIM is
powered again, on 17 March 1998 (GMT Day 076).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 25.0 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.1 degrees C. Additional battery performance data for
each battery on 07 March 1998 and 13 March 1998 are provided in Appendix D.
The clock error ranged from -10.4 to +12.3 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = none
% OUT-OF-LIMITS = 05
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% ACRIM 01
% HALOE 02
% HALOE 02
% HRDI 02
% MLS 01
% WINDII 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 01
Three (3) Anomaly Reports remained OPEN at the end of this report period.
TRANSPONDER FREQUENCIES:
Transponder A: 98/041 - 2287.497104
Transponder B: 98/042 - 2287.499330
GROUND SYSTEM OPERATION
No (0) operational support problem(s) occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/071 show the observatory data loss to be
20 hours, 42 minutes, 8.084 seconds (an increase of 0.544 seconds). This
is a 0.0365 percent data loss which equals a 99.9635 percent data capture
for the mission.
The increase of 0.544 seconds occurred on Day 067 and was due to a hit in
the data.
FOT continues to investigate the methodology of participating in the IP
Forward Link testing. The IP Forward Link testing can be completed by
immediately sending NOOP commands to the actual spacecraft (as most
projects have done), or to try to test the IP forward link configuration
with the UARS simulator. The latter option would involve having command
sent from the AP, through NASCOM, out to White Sands, back to NASCOM, and
to the simulator. At present, the FOT is unsure whether or not this
testing configuration can be accomplished.
The FOT is still unable to receive good Q-channel OBC dump data. The MSOCC
System Engineer has investigated the problem, and feels that the Q-channel
data being received at MSOCC is corrupted, and it is not the MSOCC resident
conversion devices which are corrupting the Q-channel data. Further
testing will be conducted next week.
The inability to transmit commands to the UARS Test and Training Simulator
(UTTS) was first observed on 24 February 1998. This problem cleared on 11
March 1998. The cause of the command link problem between MSOCC and the
UTTS was finally determined to be a bad NASCOM line connection.
Testing of the latest AP software (modified to support a reduced ATC load)
was successfully completed. Following the UTTS command link repair, the
ATC load and dump capabilities were tested, and all tests functioned
correctly. Once this version of the AP software is released, the ground
system will be ready to support the new PMON flight software.
Operational processing of level-0 telemetry analysis files continues in
parallel on both the UMPGAL (new MPG Alpha machine) and CDHF computer
systems. No level-0 file transfer failures (between CDHF and UMPGAL)
occurred this week.
OTHER SIGNIFICANT EVENTS
Work on the upgrade of the OBC Power Monitor (PMON) software continued.
FOT is supporting CNMOS in their efforts to make necessary software changes
(MSOCC, CMS, DCF, etc.) to support the year 2000 rollover.
UARS Memo No. 98-004, dated 09 March 1998, was issued to documents the 25
February 1998 FOT/FDF predicted attitude meeting where a decision to
execute an orbit adjust was made by the NASA Project Manager and NASA
Flight Operation Director.
UARS Memo No.98-005, dated 12 March 1998, was issued to define the planning
schedule for execution of the UARS orbit adjust in August 1998.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FOT attended a meeting with the CNMOS UARS Project Manager (John Grassel)
on 06 March 1998 to discuss Year 2000 testing requirements for the AP and
the CMS. Several actions items were assigned, with the FOT having the
responsibility of formally requesting Year 2000 planning aids from FDF.
Once the planning aids are received from FDF, CMS testing can be conducted.
This meeting was a prelude to the group meeting held on 09 March 1998.
FOT met with E.Z. Block on 10 March 1998 to discuss details involved with
the Year 2000 testing of the CMS system.
FUTURE SIGNIFICANT EVENTS
All building 23 power will be down during the weekend of 20-23 March 1998
to perform work on the building power supply. The building 23 systems will
be unavailable from about 12 pm on Friday, 20 March 1998, until about 12 pm
on 23 March 1998.
The next UARS Yaw Around Maneuver is currently scheduled for 02 April 1998.
A second building 23 power outage is scheduled from about 12 pm, 24 April
1998, until about 12 pm, 27 April 1998.
UARS Flight Operations plans to attend the Program Experience portion of
the Space Power Workshop at JPL in April 1998. UARS FOT plans on
presenting an overview of the UARS batteries and their performance.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
is targeted for 05 August 1998.
PERSONNEL CHANGES
New employee, Peter Thomas, reported for work on 09 March 1998 as an off-
line engineer.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 13 March 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
35435 066/0034 98-085 HRDI Power ON
35450 067/0037 98-086 OBC Clk Adj fr 0.00 to 20.6 msec/day
35494 069/2308 98-087 OBC Clk Adj fr 20.6 to 0.00 msec/day
35508 070/2131 98-088 SUSIM Power OFF
35508 070/2131 98-090 SUSIM RTS 12 for SUSIM
35509 070/2311 98-090 ACRIM Power OFF
35509 070/2311 98-091 HRDI Low Power Mode
35511 071/0209 98-092 MLS Power ON
35516 071/0944 98-093 OBC Clk Adj fr 0.00 to 20.6 msec/day
35527 072/0255 98-094 WINDII Power ON
35537 072/1958 98-095 HALOE Power OFF
35539 072/2319 98-096 HALOE Power ON
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 13 March 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSPBIMON Y-HI Tele Pulse Bus Current Monitor 35437
HRIFESURVTMP Y-HI HRDI IFE Survival Temp 35511
QTSISAIF Y-HI SSPP Baseplate by SISA I/F 35516,26-39
SORLBXTMPR Y-LO Logic_Box Temperature 35526-39
HATBAT Y-HI Bolometer Area Temp 35535-6
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 13 March 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
35338 059/12:13 (28 Feb 98) from 20.6 to 00.0 msec/day
35398 063/12:44 (04 Mar 98) from 00.0 to 20.6 msec/day
35425 065/07:50 (06 Mar 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 13 March 1998
07 Mar 1998 - GMT Day 066 Beta = 39.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.6 0.000 0.0
2 +67.2/ -5.6 +5.79/+4.51 +29.6/-15.2 25.6 1.026 21.9
3 +56.0/ -16.8 +4.51/+3.23 +28.8/-15.2 25.6 1.027 21.1
13 Mar 1998 - GMT Day 072 Beta = 53.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.4 0.000 0.0
2 +56.0/ -11.2 +6.44/+4.19 +26.8/-16.8 25.0 1.023 22.1
3 +84.0/ -11.2 +5.15/+2.92 +25.6/-16.8 25.0 1.024 21.6
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
APPENDIX E
UARS ANOMALY REPORTS
Week ending 13 March 1998
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-011 FOT 35533 072/1314 N/A TAC Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period